# Quantum Encrypted Control of Networked Systems

### v1

@@ Abstract
Encrypted control has been extensively studied to ensure the confidentiality of system states and control inputs for networked control systems. This paper presents a computationally efficient encrypted control framework for networked systems enabled by quantum communication. A quantum channel between sensors and actuators is used to generate identical secret keys, whose security is further enhanced through quantum key distribution. These keys enable lightweight encryption and decryption while preserving confidentiality and control accuracy. We develop a novel encryption–decryption architecture for state-feedback control of linear systems based on quantum keys, and characterize the impact of quantum-state errors on closed-loop stability. In particular, we establish the existence of a critical threshold on intrinsic quantum noise below which stability is guaranteed. In contrast to classical encrypted control schemes, which may collapse under a single key-bit error, the proposed quantum encrypted control exhibits strong robustness to key imperfections. We further adopt quantization techniques to address the scenarios with limited communication bits in practical situations, and implement privacy protection for quantum keys based on a stochastic quantizer. These results demonstrate that integrating quantum technologies into control systems in a nontrivial and principled manner, even at their current level of maturity, can yield substantial performance gains in reducing computational complexity and improving resilience to key errors while ensuring security against multiple eavesdropping sources.
Keywords: Networked systems, encrypted control, quantum entanglements
@@ 1 Introduction
    We incorporate stochastic quantization techniques to address the practical scenario of limited communication bandwidth. This approach also preserves the privacy of the quantum keys. Furthermore, we establish a trade-off between quantization error and privacy protection strength, ensuring both security and control performance of the system.
The remainder of this paper is organized as follows. Section 2 reviews networked control systems and classical symmetric and asymmetric encrypted control schemes, and refines the problem of interest. Section 3 presents the proposed $\mathsf{QEC}$ and its specific algorithmic realization. Section 4 establishes the convergence conditions under key mismatch. Section 5 applies a stochastic quantizer to $\mathsf{QEC}$, and establishes the trade-off between quantization error and the privacy protection strength of the quantum key. Section 6 reports numerical results and comparisons with existing encryption-based control methods. Finally, Section 7 concludes the paper. All proofs are collected in Appendices.
@@ 4 Stability under Intrinsic Quantum Error
Assumption 1 (Quantum Key Errors)
@@ 5.1 QEC with Quantization
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}\right)=1-(2^{w-1}y-\lfloor 2^{w-1}y\rfloor),$ (10)
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}+\frac{1}{2^{w-1}}\right)=2^{w-1}y-\lfloor 2^{w-1}y\rfloor;$
@@ 5.3 Differential Privacy of Quantum Keys
Although the introduction of quantization has introduced errors, it simultaneously provides privacy protection. Quantization helps privacy protection by converting continuous data streams into discrete values, thereby limiting information leakage. Regarding this point, existing literature has extensively investigated the privacy protection of system parameters or initial states through quantization Murguia et al. (2018); Kawano and Cao (2021); Liu et al. (2025). In this subsection, we discuss how quantization provides privacy protection to quantum keys under the notion of differential privacy. To this end, we introduce the following quantum key matrix [^1] Note that the $\beta_{i}(t),i=1,\dots,n$ are equivalent to the quantum key $q(t)$. associated with the time interval $t\in[0,T]$
@@
$\beta_{i}(t)\in\mathcal{A}_{\beta}:=\left[-2^{w_{b}-1}-\frac{\overline{x}_{i}}{\ln(1+\alpha)},-\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right)\cup\left(\frac{\overline{x}_{i}}{\ln(1+\alpha)},2^{w_{b}-1}+\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right]$
@@ 6.1 Simulation Settings
6.2 Control Performance of $\mathsf{QEC}$ Encryption with Quantum Key Errors
@@
6.3 Control Performance of $\mathsf{QEC}$ Encryption with Quantization
Next, we consider the case where the encrypted control system is affected by communication quantization. In this case, we apply Exponential-Logarithm Realization $\mathsf{QEC}$ with Quantization (13)-(16) with different bandwidth $w$ and fixed parameters $\alpha,\overline{x}_{i},\delta_{i}$. Fig. 6 illustrates the system response for different values of $w$. It can be observed in Fig. 6.(a) that the additive error induced by quantization does not vary significantly with $\mathbf{x}(t)$ but increases as the communication bandwidth decreases, which is consistent with Theorem 3.
Furthermore, we analyze how varying privacy levels affect the quantization error under a fixed communication bandwidth. To this end, we apply $\mathsf{QEC}$ defined by (13)-(16) with a sufficiently large $w$, ensuring $(0,\Delta)$-differential privacy under a fixed $\zeta$-adjacency relation for different $\Delta$. As indicated by Theorem 4, the privacy parameter $\Delta$ can be modulated by adjusting $\alpha$. The results in Fig. 6 show that stronger privacy protection (smaller $\Delta$) leads to a larger quantization error, a trend consistent with the trade-off predicted in Corollary 1.
@@ Appendix A Proof for Theorem 2
$\mathbf{x}(t+1)\mathbf{x}^{\top}(t+1)=\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{M}_{0}\mathbf{x}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top}+\mathbf{B}\mathbf{w}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{B}\mathbf{w}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top},$
@@
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{M}_{0}\mathbf{x}(t)\mathbf{B}^{\top}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}^{2}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{B}^{\top}$
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+2p\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top}+2p\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\overline{h}(p)\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top},$
@@
$\mathbf{V}(t+1)=$ $\big(\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})+\overline{h}(p)\mathbf{M}_{1}\otimes\mathbf{M}_{1}\big)\mathbf{V}(t).$
@@
$\mathbf{M}(p):=\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p\,(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})+\overline{h}(p)\,\mathbf{M}_{1}\otimes\mathbf{M}_{1},$
@@ Appendix B Proof for Theorem 3
$\leq$ $2\mathsf{E}\left(\sum_{i=1}^{n}\beta_{i}\ln(z_{i})-\sum_{i=1}^{n}[\mathbf{K}]_{i}[\mathbf{x}]_{i}\right)^{2}+2\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(\frac{g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)}{z^{1/\delta}}\right)\right)^{2}.$
@@
$\mathsf{E}\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}\right)-v\right)^{2}\leq\mathsf{E}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}-g(v)\right)^{2}\leq\frac{1}{2^{2w}},$ (30)
@@
$=$ $\sum_{i=1}^{n}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)^{2}+2\sum_{i<j}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)\mathsf{E}\left(\beta_{j}\ln\left(1+\frac{z_{j}}{z_{j}^{r}}-1\right)\right)$
@@
$\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(1+\frac{\sum_{j=0}^{w-1}2^{-j}d_{i,j}}{z^{1/\delta_{i}}}-1\right)\right)^{2}\leq\sum_{i=1}^{n}\frac{(\delta_{i}\beta_{i})^{2}}{2^{2w}}+o\!\left(\frac{1}{2^{2w}}\right),$ (36)
@@ Appendix C Proof for Theorem 4
$|\mathbb{P}(Q_{w}(v_{i}(t))\in\mathbb{S})-\mathbb{P}(Q_{w}(v_{i}^{\prime}(t))\in\mathbb{S})|\leq\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}}{2^{w-1}},\quad\forall\ \mathbb{S}\in\mathcal{F}_{0}.$ (39)

### v2

@@ Abstract
Encrypted control has been extensively studied to ensure the confidentiality of system states and control inputs for networked control systems. This paper presents a computationally efficient encrypted control framework for networked systems enabled by quantum communication. A quantum channel between sensors and actuators is used to generate identical secret keys, whose security is further enhanced through quantum key distribution. These keys enable lightweight encryption and decryption while preserving confidentiality and control accuracy. We develop a novel encryption–decryption architecture for state-feedback control of linear systems based on quantum keys, and characterize the impact of quantum state errors on closed-loop stability. In particular, we establish the existence of a critical threshold on intrinsic quantum noise below which stability is guaranteed. In contrast to classical encrypted control schemes, which may collapse under a single key-bit error, the proposed quantum encrypted control exhibits strong robustness to key imperfections. We further adopt quantization techniques to address the scenarios with limited communication bits in practical situations, and implement privacy protection for quantum keys based on a stochastic quantizer. These results demonstrate that integrating quantum technologies into control systems in a nontrivial and principled manner, even at their current level of maturity, can yield substantial performance gains in reducing computational complexity and improving resilience to key errors while ensuring security against multiple eavesdropping sources.
keywords:
Networked systems, encrypted control, quantum entanglements
thanks: A preliminary work has been submitted to the 23rd IFAC World Congress, August 23-28, 2026, BEXCO, Busan, Republic of Korea. This work was supported by the Australian Research Council under Grant DP190103615, Grant LP210200473, and Grant DP230101014, and by the Faculty of Engineering’s Breakthrough Project at The University of Sydney.
@@ 1 Introduction
    We incorporate stochastic quantization techniques to address the practical scenario of limited communication bandwidth. This approach also preserves the privacy of the quantum keys, indicating a quantum-classical confidentiality feedback loop. Furthermore, we establish a trade-off between quantization error and privacy protection strength, ensuring both security and control performance of the system.
The remainder of this paper is organized as follows. Section 2 reviews networked control systems and classical symmetric and asymmetric encrypted control schemes, and refines the problem of interest. Section 3 presents the proposed $\mathsf{QEC}$ and its specific algorithmic realization. Section 4 establishes the convergence conditions under key mismatch. Section 5 applies a stochastic quantizer to $\mathsf{QEC}$ for a fully digital implementation. Section 6 establishes the trade-off between quantization error and the privacy protection strength of the quantum key from the classical quantization. Section 7 reports numerical results and comparisons with existing encryption-based control methods. Finally, Section 8 concludes the paper. All proofs are collected in the Appendices.
Preliminary results were submitted for presentation at the 23rd IFAC World Congress. Compared to the conference version, this paper further incorporates quantization techniques, enabling the proposed algorithm to be applied in scenarios with limited communication bits. A stochastic quantizer is introduced to achieve privacy protection for the quantum keys, and the trade-off between privacy protection and quantization error is derived. Additionally, the journal version has been significantly improved, with enhanced clarity and more detailed analysis.
@@ 4 Stability under Intrinsic Quantum Error
Assumption 1 (Quantum Key Errors).
@@ 5.1 QEC with Quantization
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}\right)$ (10)
$=$ $1-(2^{w-1}y-\lfloor 2^{w-1}y\rfloor),$
$\mathbb{P}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}=\frac{\lfloor 2^{w-1}y\rfloor}{2^{w-1}}+\frac{1}{2^{w-1}}\right)$
$=$ $2^{w-1}y-\lfloor 2^{w-1}y\rfloor;$
@@
The stochastic quantizer proposed in this work is inspired by Gray and Stockham (1993). In that work, the quantizer maps each value to its two nearest quantization levels in a probabilistic manner. The proposed quantizer $Q_{w}({\cdot})$ follows the same quantization principle. However, the key difference is that $Q_{w}(\cdot)$ is defined over $w$-bit binary numbers, $\{0,1\}_{w}$, making it directly applicable in bit-constrained communication scenarios. Moreover, the function $g(\cdot)$ is initially applied to the input $v$. This is because the ciphertext to be quantized in $\mathsf{QEC}$ has an exponential form, its values within the interval $\left[\frac{1}{2},2\right]$ exhibit symmetry around $1$. Specifically, for $e^{-x}\in\left[\frac{1}{2},1\right]$, the value of $e^{x}$ lies within $\left[1,2\right]$. To exploit this symmetry, $g(v)$ in (9) leaves values of $v>1$ unchanged and maps $v\in\left[\frac{1}{2},1\right]$ to the range $\left[0,1\right]$. Subsequently, after quantizing $y=g(v)\in\left[0,2\right]$ using (10), a unified quantization error, as described in (11)-(12), can be derived for $v\in\left[\frac{1}{2},2\right]$.
[…]
@@ 5.2 Effect of Quantization Error
6 Quantum-Classical Feedback in Confidentiality
Although the introduction of quantization has introduced errors, it simultaneously provides privacy protection. Quantization helps privacy protection by converting continuous data streams into discrete values, thereby limiting information leakage. Regarding this point, existing literature has extensively investigated the privacy protection of system parameters or initial states through quantization Murguia et al. (2018); Kawano and Cao (2021); Liu et al. (2025).
In this section, we discuss how quantization provides privacy protection to quantum keys under the notion of differential privacy, completing the feedback loop of confidentiality between the control system signals and quantum keys.
@@ 6.1 Differential Privacy of Quantum Keys
We introduce the following quantum key matrix [^1] Note that the $\beta_{i}(t),i=1,\dots,n$ are equivalent to the quantum key $q(t)$. associated with the time interval $t\in[0,T]$
@@
$\beta_{i}(t)\in\mathcal{A}_{\beta}$= $\left[-2^{w_{b}-1}-\frac{\overline{x}_{i}}{\ln(1+\alpha)},-\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right)\cup$
$\left(\frac{\overline{x}_{i}}{\ln(1+\alpha)},2^{w_{b}-1}+\frac{\overline{x}_{i}}{\ln(1+\alpha)}\right]$
@@ 7.1 Simulation Settings
7.2 Control Performance of $\mathsf{QEC}$ with Quantum Key Errors
@@
7.3 Control Performance of $\mathsf{QEC}$ with Quantization
We consider the case where the encrypted control system is affected by communication quantization. In this case, we apply Exponential-Logarithm Realization $\mathsf{QEC}$ with Quantization (13)-(16) with different bandwidth $w$ and fixed parameters $\alpha,\overline{x}_{i},\delta_{i}$. Fig. 6 illustrates the system response for different values of $w$. It can be observed in Fig. 6.(a) that the additive error induced by quantization does not vary significantly with $\mathbf{x}(t)$ but increases as the communication bandwidth decreases, which is consistent with Theorem 3.
7.4 Privacy Protection of Quantum Keys via Quantization
We analyze how varying privacy levels affect the quantization error under a fixed communication bandwidth. To this end, we apply $\mathsf{QEC}$ defined by (13)-(16) with a sufficiently large $w$, ensuring $(0,\Delta)$-differential privacy under a fixed $\zeta$-adjacency relation for different $\Delta$. As indicated by Theorem 4, the privacy parameter $\Delta$ can be modulated by adjusting $\alpha$. The results in Fig. 6 show that stronger privacy protection (smaller $\Delta$) leads to a larger quantization error, a trend consistent with the trade-off predicted in Corollary 1.
@@ Appendix A Proof for Theorem 2
$\mathbf{x}(t+1)\mathbf{x}^{\top}(t+1)=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{M}_{0}\mathbf{x}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top}$
$+\mathbf{B}\mathbf{w}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathbf{B}\mathbf{w}(t)\mathbf{w}^{\top}(t)\mathbf{B}^{\top},$
@@
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{M}_{0}\mathbf{x}(t)\mathbf{B}^{\top}$
$+\mathsf{E}(\mathbf{w}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\mathsf{E}(\mathbf{w}^{2}(t)|\mathbf{x}(t))\mathbf{B}\mathbf{B}^{\top}$
$=$ $\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+2p\mathbf{M}_{0}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top}$
$+2p\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{0}+\overline{h}(p)\mathbf{M}_{1}\mathbf{x}(t)\mathbf{x}^{\top}(t)\mathbf{M}_{1}^{\top},$
@@
$\mathbf{V}(t+1)=$ $\big(\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})$
$+\overline{h}(p)\mathbf{M}_{1}\otimes\mathbf{M}_{1}\big)\mathbf{V}(t).$
@@
$\mathbf{M}(p)$= $\mathbf{M}_{0}\otimes\mathbf{M}_{0}+2p\,(\mathbf{M}_{1}\otimes\mathbf{M}_{0}+\mathbf{M}_{0}\otimes\mathbf{M}_{1})$
$+\overline{h}(p)\,\mathbf{M}_{1}\otimes\mathbf{M}_{1},$
@@ Appendix B Proof for Theorem 3
$\leq$ $2\mathsf{E}\left(\sum_{i=1}^{n}\beta_{i}\ln(z_{i})-\sum_{i=1}^{n}[\mathbf{K}]_{i}[\mathbf{x}]_{i}\right)^{2}$
$+2\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(\frac{g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}d_{i,j}\right)}{z^{1/\delta}}\right)\right)^{2}.$
@@
$\mathsf{E}\left(g^{-1}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}\right)-v\right)^{2}$ $\leq\mathsf{E}\left(\sum_{j=0}^{w-1}2^{-j}a_{j}-g(v)\right)^{2}$ (30)
$\leq\frac{1}{2^{2w}},$
@@
$=$ $\sum_{i=1}^{n}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)^{2}$
$+2\sum_{i<j}\mathsf{E}\left(\beta_{i}\ln\left(1+\frac{z_{i}}{z_{i}^{r}}-1\right)\right)\mathsf{E}\left(\beta_{j}\ln\left(1+\frac{z_{j}}{z_{j}^{r}}-1\right)\right)$
@@
$\mathsf{E}\left(\sum_{i=1}^{n}\delta_{i}\beta_{i}\ln\!\left(1+\frac{\sum_{j=0}^{w-1}2^{-j}d_{i,j}}{z^{1/\delta_{i}}}-1\right)\right)^{2}$ (36)
$\leq$ $\sum_{i=1}^{n}\frac{(\delta_{i}\beta_{i})^{2}}{2^{2w}}+o\!\left(\frac{1}{2^{2w}}\right),$
@@ Appendix C Proof for Theorem 4
$|\mathbb{P}(Q_{w}(v_{i}(t))\in\mathbb{S})-\mathbb{P}(Q_{w}(v_{i}^{\prime}(t))\in\mathbb{S})|$ (39)
$\leq$ $\zeta\frac{(1+\alpha)\ln^{2}(1+\alpha)}{\overline{x}_{i}}{2^{w-1}},\quad\forall\ \mathbb{S}\in\mathcal{F}_{0}.$
@@ Appendix D Proof for Corollary 1
Acknowledgment
The authors thank Elizabeth Ratnam (Monash University) and Farhad Farokhi (The University of Melbourne) for insightful discussions that inspired the early ideas of this paper.